hallm:

naomi deposited money into a savings account that is compounded quarterly at an interest rate of 6%. She thinks this quarterly rate is the same as a mounthly intrest rate of 1.25%. Is naomi right? If so, eplains why and show how you got that answer. Is naomi inncorrect? If so explain the mistake in her reasoning, give the correct monthly interest rate, and show how you got your answer.

jhonyy9:

any idea ? do you know what mean quarterly rate ?

jhonyy9:

you can checking here : https://www.thefreedictionary.com/quarterly

jhonyy9:

hallm are you here ? please collaborate - i like help you

jhonyy9:

quarter of a year = 12 months divide 4 = 3 month so from this result that the quarterly rate of 6% mean a rate of 3 month yes ?

hallm:

ohh im sorry i didn't noticed that someone was on here i was doing something im here now though

jhonyy9:

ok so what is your idea now ? can you continue it ?

hallm:

i think so if i need more help i can always message you to see if your on to help if that is ok with you

jhonyy9:

ok - but please write mea here how you solve this your exercise ?

hallm:

ok i will

jhonyy9:

wait it

hallm:

??

jhonyy9:

i wait your work - please write here how you solve it

jhonyy9:

you ve said that you understand what i ve wrote above

hallm:

im a little confused at what your saying

jhonyy9:

so you know that quarterly mean 12/4 = 3 months Naomi thinks that this quarterly rate is the same as a monthly rate of 1,25% but you know that a quarterly rate mean a rate of 3 months what is given that is 6% in this case you can writing, using the rule of 3 simple ,so a rate of 3 months equal 6% a rate of 1 month equal x% x = 1*6/3 = 6/3 = 2% so from this result that the Noemi's answer not is right

jhonyy9:

do you understand it now ?

hallm:

yes i do

jhonyy9:

@ThisGirlPretty hallm sure ?